Consider the following Decision alternative for the Raman Pahwa, he wants to invest in


stocks, and thought about two situations about tomorrow’s market condition. The figures


(in INR) in the following table exhibit profit per unit of stock-investment.


payoff table:


Favorable market Unfavorable market


Lakshmi pvt ltd 55 26


Mehta Groups of industries 43 38


Surya 29 43


LT energy 15 51


1. Draw the decision tree


2. If we assign the following probabilities to the states of nature, then determine the


EMV decision.


P(s1) = .4 P(s2) = .1 P(s3) = .3 P(s4) = .2

1
Expert's answer
2022-02-23T13:07:03-0500


The above is the decision tree.

EMV = multiplying the value of each possible outcome (impact) by its likelihood of occurrence (probability) and then adding the results

(4*2) + (4*3) = 8+12= 20